3. Susans widget factory produces as many as 1,200 widgets daily, and sells them for $26 each. If the rent on the plant and equipment is $13,000 per month, the raw material for each widget is $7.50 and workers are paid $19.00 per hour and one (1) worker can produce (30) widgets per hour - then: What is the breakeven point? Point at which costs and revenues are equal. What is the maximum profit potential of this plant? What happens to the breakeven point and profit potential if a worker productivity increases to 31 widgets per hour? a. b. c.
Answer #1

Answer a : Break even point = Fixed cost / Price - Variable cost

Break Evan point = 13000/ 26-(7.50+0.63)

Break even point =13000/ 17.87=728 units

It is the point where total cost is exactly equal to total revenue.

Total Revenue = Price × Quantity= $26×728=$18,928

Total cost =13000+5460+458.60=$18928

B : Maximum profit potential of the firm

Total revenue = 1200*$26=$31200

Total cost =$13000+$9000+$756= $22,756

Maximum profit = $8444

C : It means variable cost has been decreased as

Break even point = 13000/ 26-(7.50+0.61)=727 units

Total revenue =727*$26=$18902

Total cost = 13000+5452+ 440=$18902